OpenStudy (anonymous):

y+2=-3/5(x+10) in standard form

OpenStudy (anonymous):

distribute your 3/5 to (x+10), then subtract the 2 to the other side.

OpenStudy (anonymous):

i dont understand still

OpenStudy (anonymous):

do you know what the distribution property is? a(b+c) = ab+ac

OpenStudy (anonymous):

so it is 3/5x +6-y=2

OpenStudy (anonymous):

y=-3/5x

OpenStudy (anonymous):

wait... woops.. error...y=-3/5x-8

Directrix (directrix):

The equation is not yet written in standard form as dictated by the question. Standard form: Ax + By = C. So, we have: y=-3/5x-8 Multiplying by 5 yields: 5y = -3x - 40 3x + 5y = 40 --> is the given equation in standard form http://jwilson.coe.uga.edu/emt668/emat6680.2002/jackson/chapter%205%20lesson%20plan/day6.html